# Alignment and Adversarial Robustness: Are More Human-Like Models More Secure?

### v1

@@ 5 Related Work
There has been substantial progress on bridging the representational differences between humans and machine learning models over the last few years. Geirhos et al. (2021) shows many of the high-performance models match or in many cases exceed human feedforward performance on most of the OOD datasets studied. New models have also been introduced to promote both alignment and robustness. For example, Dapello et al. (2020) designs a new block for CNNs called the VOne block, which simulates V1 area processing. This work found that incorporating the VOne block into ResNet models increased robustness to both white box adversarial examples and common corruptions without sacrificing clean performance on ImageNet. Li et al. (2019) introduced a technique for regularizing machine learning models based on human neural readings and found that the resultant regularized models were more robust and human-aligned.
Subramanian et al. (2023) shows that the property difference of the spatial frequency channel between humans and neural networks explains both shape bias and adversarial robustness of networks. Models with higher levels of human alignment have also been shown to be more robust to distribution shifts and ImageNet-A data Sucholutsky & Griffiths (2023). Additionally, it has been shown that models tend to prioritize texture information over shape information Geirhos et al. (2019); Hermann et al. (2020) and that this bias extends to real-data decisions and is one of the major causes for vulnerability to natural adversarial samples Hoak et al. (2024); Hoak & McDaniel (2024).
@@ 6 Conclusions
In this work, we find that, perhaps surprisingly, representational alignment and adversarial robustness in vision systems are not always correlated. However, we do observe that certain individual benchmarks serve as strong indicators of robust accuracy, particularly those that assess a model’s preference for texture information over shape. From this, we hope to encourage future work to leverage insights found in both areas to build more secure and aligned vision systems.
Acknowledgments
@@
Appendix A Appendix
Evaluated Models
grcnn vgg 19 vgg 16
alexnet CORnet-S ReAlnet10
ReAlnet09 ReAlnet08 ReAlnet07
ReAlnet05 ReAlnet04 ReAlnet03
ReAlnet02 grcnn 109 ReAlnet01
BiT-S-R50x1 AlexNet SIN BiT-S-R50x3
BiT-S-R101x1 densenet-121 densenet-169
resnet50-sup resnet50-SIN inception v1
nasnet large BiT-S-R152x2 densenet-201
BiT-S-R152x4 inception v4 inception v3
BiT-S-R101x3 nasnet mobile artResNet18 1
regnet y 400mf antialiased-r50 resnet50-barlow
imagenet l2 3 0 efficientnet b0 resnet50-SIN IN
effnetb1 272x240 resnet-50-robust Res2Net50 26w 4s
resnet50 tutorial convnext tiny sup resnet50-SIN IN IN
shufflenet v2 x1 0 AT efficientnet-b2 tv efficientnet-b1
resnet SIN IN FT IN resnet50-vicregl0p9 ViT L 32 imagenet1k
alexnet ks torevert antialias-resnet152 resnet50-vicregl0p75
resnext101 32x8d wsl resnet-152 v2 pytorch resnext101 32x16d wsl
resnext101 32x32d wsl resnext101 32x48d wsl resnet50 imagenet full
resnet34 imagenet full resnet18 imagenet full focalnet tiny lrf in1k
resnet50 robust l2 eps1 AdvProp efficientnet-b4 resnet50 robust l2 eps3
AdvProp efficientnet-b6 resnet152 imagenet full resnet101 imagenet full
AdvProp efficientnet-b8 AdvProp efficientnet-b7 AdvProp efficientnet-b2
convnext femto ols:d1 in1k antialiased-rnext101 32x8d resnet50 imagenet 1 seed-0
resnet50 imagenet 10 seed-0 cv 18 dagger 408 pretrained convnext tiny:in12k ft in1k
deit small imagenet 1 seed-0 resnet50 imagenet 100 seed-0 efficientnet b0 imagenet full
efficientnet b2 imagenet full efficientnet b1 imagenet full deit small imagenet 10 seed-0
deit base imagenet full seed-0 deit small imagenet 100 seed-0 deit large imagenet full seed-0
convnext large:fb in22k ft in1k deit small imagenet full seed-0 convnext small imagenet 1 seed-0
convnext xlarge:fb in22k ft in1k convnext small imagenet 10 seed-0 convnext base imagenet full seed-0
convnext tiny imagenet full seed-0 convnext small imagenet 100 seed-0 convnext small imagenet full seed-0
convnext large imagenet full seed-0 vit base patch16 clip 224:openai ft in1k vit large patch14 clip 224:openai ft in1k
vit large patch14 clip 224:laion2b ft in1k convnext xxlarge:clip laion2b soup ft in1k vit tiny r s16 p8 384:augreg in21k ft in1k
vit large patch14 clip 336:laion2b ft in1k vit relpos base patch16 clsgap 224:sw in1k effnetb1 cutmix augmix sam e1 5avg 424x377
vit relpos base patch32 plus rpn 256:sw in1k convnext base:clip laiona augreg ft in1k 384 vit base patch16 clip 224:openai ft in12k in1k
swin small patch4 window7 224:ms in22k ft in1k vit huge patch14 clip 224:laion2b ft in12k in1k vit large patch14 clip 224:openai ft in12k in1k
vit large patch14 clip 336:openai ft in12k in1k vit huge patch14 clip 336:laion2b ft in12k in1k vit large patch14 clip 224:laion2b ft in12k in1k
convnext large mlp:clip laion2b augreg ft in1k 384 resnet50 finetune cutmix AVGe2e3 robust linf8255 e0 247x234 effnetb1 cutmixpatch augmix robust32 avge4e7 manylayers 324x288
effnetb1 cutmixpatch SAM robust32 avge6e8e9e10 manylayers 324x288

### v2

@@ 5 Related Work
There has been substantial progress in bridging the representational differences between humans and machine learning models over the last few years. Geirhos et al. [15] show that many of the high-performance models match or exceed human feedforward performance on OOD datasets. Several works suggest that more human-aligned model architectures may also be more robust. Dapello et al. [10] design the biologically inspired VOne block to simulate V1 processing, improving robustness to adversarial and common corruptions. Li et al. [25] regularize models based on neural recordings from mice to increase robustness and alignment.
Furthermore, Subramanian et al. [35] show that differences in spatial frequency processing explain both shape bias and adversarial vulnerability. Models with higher human alignment show improved performance on datasets like ImageNet-A [36]. Other works highlight that models tend to rely on texture over shape [13, 18], a bias that increases susceptibility to natural adversarial samples [20, 19].
Our work expands on these findings with a large-scale empirical study across diverse architectures and benchmarks. Unlike prior work that focuses on specific models or alignment methods, we systematically analyze how different forms of alignment relate to adversarial robustness, discovering meaningful connections between model perception and security.
@@ 6 Conclusions
In this work, we investigated the relationship between representational alignment and adversarial robustness in vision models. While robustness showed only a weak correlation with average alignment scores, we found that specific alignment benchmarks, particularly those measuring texture selectivity, were strongly predictive of robust accuracy. Additionally, we observed that robustness is well-structured in alignment space: models with similar alignment profiles tend to exhibit similar levels of robustness, and highly robust models occupy a distinct region of this space. These findings suggest that not all alignment is equally beneficial for security and that targeted alignment along specific dimensions is a promising strategy for improving robustness. We hope this work motivates future research at the intersection of alignment and security, with the goal of developing vision systems that are both perceptually grounded and resilient to adversarial threats.
{ack}